OpenStudy (anonymous):

solve the equation (8/7)^3=343/512

OpenStudy (alexandervonhumboldt2):

\[(\frac{ 8 }{ 7 })^3=\frac{ 343 }{ 512 }\] this is not an equation

OpenStudy (alexandervonhumboldt2):

this is a false statement

OpenStudy (alexandervonhumboldt2):

because (8/7)^3 is not 343/512

OpenStudy (anonymous):

the answers provided are A. {-1/3} B. {1/3} C. {-3} D. {3}

OpenStudy (owlcoffee):

Please recheck if you have well written the problem.

OpenStudy (anonymous):

yup,agree with @AlexandervonHumboldt2 this is not an equation but it is a false statement pls check ur question again

OpenStudy (anonymous):

(8/7)^x=343/512

OpenStudy (alexandervonhumboldt2):

this is better

OpenStudy (alexandervonhumboldt2):

\[(\frac{ 8 }{ 7 })^x=\frac{ 343 }{ 512 }\]

OpenStudy (alexandervonhumboldt2):

look: 7^3=343 8^3=512

OpenStudy (alexandervonhumboldt2):

a^{-b}=1/(a^b)

OpenStudy (anonymous):

\[(\frac{ 8 }{ 7 })^x=\frac{ 7^3 }{ 8^3 }\]

OpenStudy (alexandervonhumboldt2):

oh yeah marc wrote this correctly

OpenStudy (anonymous):

thnx @AlexandervonHumboldt2 :)

OpenStudy (alexandervonhumboldt2):

your exponent should be negative

OpenStudy (alexandervonhumboldt2):

Negative Power Rule: x^{-a}=1/(x^a)

OpenStudy (anonymous):

:D I actually have no idea. I was asking for my friends little sister. Thanks a bunch.

OpenStudy (anonymous):

\[Example,(\frac{ 1 }{ 2 })^t=(\frac{ 2 }{ 1 })^{-t}\]

OpenStudy (alexandervonhumboldt2):

\[(8/7)^{x}=1/\frac{ 512 }{ 343 }\]

OpenStudy (anonymous):

\[(\frac{ 8 }{ 7 })^x=(\frac{ 1 }{ \frac{ 343 }{ 512 } })^{-1}\]

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

Thanks she understands the problem now.

OpenStudy (anonymous):

yw